VEGAS Team Meeting

July 08, 2013

Attendees

CALTECH, UC-B, & Columbia

Dan, Glenn, Mark,

NRAO:

Marty, Anish. John, Jason, Mark

Agenda:

  • Testing
Test results

Meeting Notes:

Test results ADC Testing - 1/2 Nyquist and nyquist spurs are seen and compensation code may be written. Don't know if that code exists somewhere (outside of project).
  • next will be crosstalk tests - req. cross-coupling test generator
  • Connect to ROACH 4 ; power level ~ -10dBm (Jason)

PFB tets ~ -6dB vs. target of -3dB. May be a calculation issue and plots for the team later this week. Is 3dB a requirement? Side lobe level is a better attribute to use for acceptance. Anish to evaluate, additional discussion on the VEGAS internal teleconference 7/8/13

Computers - Tracking info will be sent by UC-B scheduled for Friday delivery

CPUs sent from UC-B were received in GB

Valon control software is needed to get rid of serial port connections to a couple of ROACHs.

Mode 1 data is not yet available for nodding

Issues (All modes)
  1. Number of integrations issue
  2. Normalization of the spectrum
  3. FITS and file writing optimization for highest speed data rates (header update)

Analog tests by Galen

Next Teleconference: July 15, 2013 .

Long-term action Items
  1. Ray to make VEGAS simulator.
  2. A BOF for mode 13, or some other mode that uses GPUs.
    • (04-Feb) Not available yet
    • (11-Feb) 16 input version compiling now for subsequent test
    • (04-Mar) Should be ready to test in GB ~7-Mar
      • Need to refine what tests are done in Berkeley v. GB
        • Noise + tone tests in Berkeley
        • Replicate in GB (one Roach)
        • Build-out to eight Roach boards
        • Test on GBT
  3. A BOF with working switching signal logic.
    • (04-Feb) designed but still troubleshooting switching table loading issues
    • (11-Feb) Added to design for Mode #1
    • (11-Feb) Add sfp shift into manager
  4. A test and decision on what computers to buy.
    • (04-Feb) Not yet determined - Jayanth should go ahead with his tests to complete selection
    • (04-Feb) Needs HPC11.
  5. All Banks not starting at the same time.
    • (04-Feb) (above)
    • DONE (04-Mar) Resolved
  6. It will be good if we get the full spectral dynamic range corresponding to the 8 bit ADC in the next version of mode 1.
    • (04-Feb) Conduct initial tests at Berkeley, to see if there are any limiting issues for dynamic range and optimize
    • (04-Feb) what is/are the requirement(s) w.r.t full scale sin wave?
    • (04-Feb) is this a new requirement for the manager (yes)? is a parameter required to select between options? (parametrized shifting)
    • (04-Feb) what are requirements for shift parameter? depends on result of studying spectra...
    • (25-Feb)
  7. KFPA observation with mode 13.
    • (04-Feb) awaits mode 13
  8. Manager/VEGAS HPC code that will let two instances of the program be run simultaneously.
    • (04-Feb) not available, but not required for GC obs.
  9. Remember to do the RFI test of the rack.
    • (04-Feb) Not yet scheduled
  10. Roach to Valon serial port connection.
    • (04-Feb) Not yet implemented; not req'd for GC obs
  11. The I2C bus software tested with the manager and Galen's modules.
    • (04-Feb) Post GC Obs
  12. Add software version information into FITS file
    • (11-Feb) After GC obs
  13. Confine UDP and packets out of the subnet within the VEGAS switch

Other items:

-- MartyBloss - 2013-07-08
Topic revision: r2 - 2013-07-08, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback